OpenStudy (el_arrow):

find the partial fraction decomposition for the following rational expression

OpenStudy (el_arrow):

\[\int\limits_{?}^{?} 2x^4-8x^2+5x-2/x^3-4x\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you have to divide first

OpenStudy (el_arrow):

do i factor out an x at the bottom

OpenStudy (misty1212):

no not yet

OpenStudy (misty1212):

you cannot do the partial fractions with the degree of the numerator larger than the degree of the denominator you have to do the long division first actually divide in other words

OpenStudy (el_arrow):

okay let see

OpenStudy (misty1212):

i would use technology to divide, long division is a pain

OpenStudy (el_arrow):

yeah i got my calculator right beside me

OpenStudy (misty1212):

lol why use a calculator? you are on a computer right? http://www.wolframalpha.com/input/?i=+%282x%5E4-8x%5E2%2B5x-2%29%2F%28x%5E3-4x%29

OpenStudy (misty1212):

of course wolfram will give the the final answer as well, but if you want the quotient and the remainder, it is \[2x+\frac{5x-2}{x^3-4x}\]

OpenStudy (el_arrow):

i dont have gopro though

OpenStudy (el_arrow):

i dont think i am suppose to use long division

OpenStudy (misty1212):

i don't know what go pro is you have to divide there is no choice

OpenStudy (misty1212):

the partial fractions is for the part \[\frac{5x-2}{x^3-4x}\] the integral of the first part is \(x^2\)

OpenStudy (el_arrow):

its asking to get the final answer not the quotient and remainder

OpenStudy (misty1212):

i guess i am not being clear in order to use partial fractions to solve this integral, you have to divide the integrand to find the quotient and remainder the partial fractions is only for the remainder the quotient is \(2x\) and when you integrate that you get \(x^2\) the integral you don't know is \[\int \frac{5x-2}{x^3-4x}dx\] and that you solve using partial fractions